OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

y=-3x+7 To become perpendicular: Has to be negative reciprocal y=1/3x, can eliminate c and d

OpenStudy (anonymous):

y=1/3x+b -1=2+b b=-3, y=1/3x-3

OpenStudy (anonymous):

Is it b or a?

OpenStudy (anonymous):

a is the answer